# Generalized Multi-Objective Reinforcement Learning with Envelope Updates in URLLC-enabled Vehicular Networks

### v1

@@ VI Conclusion
Addressing the challenge of learning optimized policies across varied preferences, we develop an envelope MORL solution. This approach enables effective navigation across preference spaces, generating tailored policies. Our algorithm leverages the contraction properties of the optimality operator governing a generalized Bellman equation and optimizes the convex envelope of multi-objective Q-values. Hindsight experience replay and homotopy optimization aid in manageable learning across diverse preferences. Additionally, we construct a novel simulation testbed, ”RF-THz-Highway-Env,” based on ”highway-env,” emulating a multi-band wireless network-enabled VNet. Numerical results demonstrate the superiority of our proposed solution over weighted sum-based MORL solutions with DQN, showcasing improvements of 12.7%, 18.9%, and 12.3% on average transportation reward, average communication reward, and average HO rate, respectively.

### v2

@@ V-C5 Pareto Front Analysis
V-C6 Training and Evaluation Time Complexity Per Step
We trained each algorithm for 4000 episodes, with up to $T_{hl}$ steps per episode, resulting in a total duration per step of approximately 0.643 seconds. The breakdown of time per step during training and evaluation is presented in Table V. The reward objectives are modified as follows:
  • Telecommunication Only This scenario considers only the telecommunication reward in the MDP design. The telecommunication reward has no impact on the training process.
  • Transportation Only This scenario focuses solely on the transportation reward, with the telecommunication reward not influencing the training.
  • Telecommunication+Transportation This scenario incorporates both the transportation and Telecommunication rewards jointly during the training process.
Instance Training (s) Evaluation (s)
Transportation Telecommunication + Transportation Transportation Telecommunication + Transportation
I-(20,30,20,20) 0.0743 0.0754 0.0643 0.0654
I-(25,35,20,20) 0.0724 0.0731 0.0614 0.0613
I-(20,30,10,20) 0.0764 0.0784 0.0637 0.0648
I-(20,30,20,50) 0.0826 0.0845 0.0667 0.0672
I-(30,40,20,20) 0.0858 0.0903 0.0725 0.0736
Table V: Comparison of Training and Evaluation Times for Different Instances (in seconds).
It is important to note that the time consumption of Telecommunication+Transportation is reasonable and comparable to the system policy update 0.067 s (15 Hz) [24]. In [30, 24, 41], the authors considered AV motion updates at a rate of 15 Hz, i.e., 0.067s or 67 ms. Note that, the update frequency is not dependent on the desired velocity. On the other hand, according to uRLLC (5), the transmission latency is given by $D_{t}=\frac{L_{B}}{W}$, where $W=W_{T}$ for TBSs and $W=W_{R}$ for RBSs, thus $D_{t}=5\times 10^{-8}$s for TBSs and $D_{t}=4\times 10^{-7}$s for RBSs [47, 48], which are relatively smaller than $0.0667$s. As such, the telecommunication actions can update at the same time scale as the AV motion action updates. Thus, updating and synchronizing telecommunication and AV motion updates in environments such as highway-env [24] or the proposed RF-THz-highway-env is reasonable.
V-C7 Simultaneously Multi Objective Optimization Analysis
Figure 8: MO-Envelope-DDQN performance comparison with (1) Telecommunication optimization only, (2) Transportation optimization only, and (3) Joint Telecommunication + Transportation optimization.
As depicted in Figure 8, we consider the following three scenarios to examine the impact of optimizing different objectives. we found that considering both telecommunication and transportation objectives together performs well compared to relying on Telecommunication Only and Transportation Only alone.
V-C8 MORL Benchmarks Training Time Comparison
Instance MO-Q[44] MO-DQN [1] MO-Dueling-DDQN [11] MO-PPO [45] EUPG [46] MO-Envelope-DDQN
I-(20,30,20,20) 30:36 31:10 30:45 31:50 32:05 30:15
I-(25,35,20,20) 27:24 27:22 28:59 28:41 28:17 26:25
I-(20,30,10,20) 30:11 30:53 30:32 31:22 31:57 30:06
I-(20,30,20,50) 32:00 32:20 31:50 33:00 33:15 31:45
I-(30,40,20,20) 21:45 22:08 21:33 22:52 23:35 21:17
Table VI: Comparison of Training Times for Different MORL Methods. The times reported in the tables are formatted as MM:SS, where MM represents the number of minutes and SS represents the number of seconds.
We trained each algorithm for 4000 episodes on the test instances given in Table III and recorded the total training time, as summarized in Table VI. Our proposed MO-Envelope-DDQN demonstrates competitive training times while offering significant improvements in multi-objective optimization performance. Our results in Table VI show that the training time is mostly comparable to the baselines, if not significantly less.
@@ VI Conclusion
Addressing the challenge of learning optimized policies across varied preferences, we develop an envelope MORL solution. This approach enables effective navigation across preference spaces, generating tailored policies. Our algorithm leverages the contraction properties of the optimality operator governing a generalized Bellman equation and optimizes the convex envelope of multi-objective Q-values. Hindsight experience replay and homotopy optimization aid in manageable learning across diverse preferences. Additionally, we construct a novel simulation testbed, ”RF-THz-Highway-Env,” based on ”highway-env,” emulating a multi-band wireless network-enabled VNet. Numerical results demonstrate the superiority of our proposed solution over weighted sum-based MORL solutions with DQN, showcasing improvements of 12.7%, 18.9%, and 12.3% on average transportation reward, average telecommunication reward, and average HO rate, respectively.
Future research could extend the generalization of the MORL model, which is necessary to adapt to the dynamics of the traffic flow.